OpenStudy (anonymous):

How many terms at least of the A.P 1,4,7,10,.... are needed to get a sum greater than 590? . . ANSWER: n=21

hartnn (hartnn):

know the formula for sum of AP ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

S=n/2(2a+(n-1)d) The Sum should be greater than 590 590<n/2(2a+(n-1)d) Here a=1,d=3 Substitute and find n...

OpenStudy (anonymous):

Sn=n/2[2a+(n-1)d]

hartnn (hartnn):

yes, that ^

hartnn (hartnn):

where r u stuck ?>

OpenStudy (anonymous):

i cant understand the question

OpenStudy (anonymous):

how will i do it to get sum greater than 590

hartnn (hartnn):

plug in S = 590 , a=1,d= 3 in the formula.

hartnn (hartnn):

sum greater than 590 --->Sn > 590

hartnn (hartnn):

n/2[2a+(n-1)d] > 590

OpenStudy (anonymous):

Then u will get fraction in n.........might be be....somethin 20.345..like tht.... 21 will be the rqd n...since sum of 20 terms is less thn 590...

OpenStudy (anonymous):

Understood?

hartnn (hartnn):

what are you waiting for? start substituting....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

n/2[2a+(n-1)d] > 590 what should i do with this greater sign?

hartnn (hartnn):

keep it as it is... put, a, d values, and you'll get a quadratic in n factorize it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

n/2(2+3n-3)>590

OpenStudy (anonymous):

is it right

hartnn (hartnn):

yes, go ahead and bring it in the form ax^2+bx+c >0

OpenStudy (anonymous):

2n+3n^2-3n/2 >590 (can i cross multiply 2)

hartnn (hartnn):

yes, as 2 is positive, you can cross multiply without any sign change.

OpenStudy (anonymous):

3n^2-n-1180>0

hartnn (hartnn):

and don't you get n+ 3n^2/2 -3n/2 >590

hartnn (hartnn):

n/2(2+3n-3)>590 n (3n-1) > 590*2 3n^2 -n - 1180 >0 yes, thats correct.

OpenStudy (anonymous):

can the middle term break?

OpenStudy (anonymous):

or i have to use quadratic formula?

OpenStudy (anonymous):

??

hartnn (hartnn):

apparently, yes. but its hard to see. i would suggest quadratic formula only..

hartnn (hartnn):

see, whether u get (x-20) (3x+59) > 0

OpenStudy (anonymous):

so can u tell me how is the middle term breaking?

OpenStudy (anonymous):

ok got it

hartnn (hartnn):

1180*3 = 59*60

OpenStudy (anonymous):

i solved it by quadratic

hartnn (hartnn):

you must have got one positive(20 )and one negative solution ? igone negative.

hartnn (hartnn):

*ignore

OpenStudy (anonymous):

n=20 and n=-59/3

hartnn (hartnn):

yes, so n-20 is a factor. n-20 > 0 ---> n>20 ---->n=21

hartnn (hartnn):

got it ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^